### v2

# A Comprehensive Survey of Scientific Large Language Models and Their Applications in Scientific Discovery

Abstract
In many scientific fields, large language models (LLMs) have revolutionized the way text and other modalities of data (e.g., molecules and proteins) are handled, achieving superior performance in various applications and augmenting the scientific discovery process. Nevertheless, previous surveys on scientific LLMs often concentrate on one or two fields or a single modality. In this paper, we aim to provide a more holistic view of the research landscape by unveiling cross-field and cross-modal connections between scientific LLMs regarding their architectures and pre-training techniques. To this end, we comprehensively survey over 250 scientific LLMs, discuss their commonalities and differences, as well as summarize pre-training datasets and evaluation tasks for each field and modality. Moreover, we investigate how LLMs have been deployed to benefit scientific discovery. Resources related to this survey are available at https://github.com/yuzhimanhua/Awesome-Scientific-Language-Models. * Equal contribution
A Comprehensive Survey of Scientific Large Language Models and
Their Applications in Scientific Discovery
[…]
1 Introduction
The emergence of large language models (LLMs) Zhao et al. (2023c) brings a new paradigm to natural language processing (NLP) by replacing specialized models designed for each task with unified models that are reasonably effective for a wide spectrum of problems. In the scientific domain, such a paradigm not only reshapes people’s strategies to handle tasks related to natural language (e.g., scientific papers, medical records, and climate reports) but also inspires analogous ideas to deal with other types of data (e.g., molecules, proteins, tables, and metadata). In addition to understanding existing scientific data, LLMs have shown their potential to accelerate scientific discovery Wang et al. (2023c); Zhang et al. (2023e); Wang et al. (2024d) through generation, planning, etc.
Figure 1: Three major types of scientific LLM pre-training techniques. (Column 1): Pre-training encoder LLMs with sequentialized scientific data (e.g., text, academic graphs, molecules, biological sequences) via masked language modeling. (Column 2): Pre-training (encoder-)decoder LLMs with sequentialized scientific data (e.g., text, tables, crystals, images) via next token prediction (possibly with instruction tuning). (Column 3): Mapping text and relevant sequences/graphs/images closer in the latent space via contrastive learning.
Given the broad and profound impact of LLMs in various scientific fields across diverse modalities, it becomes necessary to comprehensively review related work in this direction. However, existing scientific LLM surveys typically focus on either one or two fields (e.g., biomedicine Wang et al. (2023a); He et al. (2024); Pei et al. (2024); Zhang et al. (2024d) and chemistry Xia et al. (2023); Pei et al. (2024); Zhang et al. (2024d)) or one modality (e.g., text Ho et al. (2024)) only. In fact, if we take a holistic view of the research landscape, we can observe similar and interrelated techniques used to develop LLMs for different fields and modalities.
Figure 1 depicts three major types of scientific LLM pre-training strategies (i.e., Columns 1 to 3), for each of which we give 4 examples (i.e., Types a to d). In Column 1, following BERT Devlin et al. (2019) and RoBERTa Liu et al. (2019), existing studies use masked language modeling (MLM) to pre-train encoder language models. Here, the input can be naturally sequential (e.g., papers in each field, protein/DNA/RNA sequences in the FASTA format Lipman and Pearson (1985)) or artificially linearized (e.g., molecules in the SMILES format Weininger (1988), sequences of venue/author/reference nodes in citation graphs). In Column 2, inspired by GPT Brown et al. (2020) and LLaMA Touvron et al. (2023a), previous studies adopt next token prediction to pre-train (encoder-)decoder language models, some of which further adopt instruction tuning and preference optimization Ouyang et al. (2022). Other than plain text input (e.g., question-answer pairs from knowledge bases or exams), we see more ways to sequentialize complex scientific data, such as flattening table cells and using particle coordinates to describe crystals. Even for images, there are studies in both mathematics Gao et al. (2023) and biomedicine Li et al. (2023a) that exploit a vision encoder to project an image onto several visual tokens and prepend them to text tokens as linearized LLM input. In Column 3, following DPR Karpukhin et al. (2020) and CLIP Radford et al. (2021), two encoders are pre-trained to map relevant data pairs closer in the latent space via contrastive learning. When both modalities are sequential (e.g., text-text or text-protein), the model is built upon two LLM encoders. When we prefer to keep the non-sequential nature of one modality (e.g., molecular graphs Edwards et al. (2021), chest X-rays Zhang et al. (2022), and aerial views Yan et al. (2024)), the corresponding graph or image encoder can be employed. To summarize, a cross-field, cross-modal survey will more accurately draw the connections between different scientific LLMs, demonstrate their commonalities, and potentially guide their future designs.
Contributions. In this paper, motivated by the discussions above, we systematically survey over 250 scientific LLMs encompassing various fields (e.g., general science, mathematics, physics, chemistry, materials science, biology, medicine, and geoscience), modalities (e.g., language, graph, vision, table, molecule, protein, genome, and climate time series), and sizes (from $\sim$100M to $\sim$100B parameters). For each field/modality, we investigate commonly adopted pre-training datasets, model architectures, and evaluation tasks of scientific LLMs. Following our motivation, when we discuss model architectures in detail, we link them back to Figure 1 to build cross-field cross-modal connections. Moreover, we provide a structured summary of these scientific LLMs in Table A1-Table A6 (Appendix A). Furthermore, for different fields, we introduce how LLMs have been deployed to benefit science by augmenting different aspects and stages of the scientific discovery process, such as hypothesis generation, theorem proving, experiment design, drug discovery, and weather forecasting.
2 LLMs in General Science (Table A1)
2.1 Language
The most commonly used pre-training corpora for scientific LLMs are research papers from bibliographic databases, such as AMiner Tang et al. (2008), Microsoft Academic Graph (MAG) Sinha et al. (2015), and Semantic Scholar Ammar et al. (2018). Some of these sources (e.g., S2ORC Lo et al. (2020)) contain full-text information of papers, while the others have titles and abstracts only.
The evolution of scientific LLMs bears similarity to that of general-domain LLMs. Specifically, pioneering models utilize paper text in a self-supervised manner during pre-training, aiming to acquire scientific knowledge from large-scale unlabeled corpora. For example, masked language modeling (MLM) is the default pre-training task for scientific LLMs with a BERT backbone (Type 1.a in Figure 1, e.g., SciBERT Beltagy et al. (2019)); next token prediction is widely used for GPT-based scientific LLMs (Type 2.a in Figure 1, e.g., SciGPT Luu et al. (2021)). More recently, inspired by the fact that LLMs can be trained to follow natural language instructions Wei et al. (2022a); Ouyang et al. (2022), researchers have put more effort into tuning LLMs with instructions to solve complex scientific problems (Type 2.a, e.g., Galactica Taylor et al. (2022) and SciGLM Zhang et al. (2024a)). The instruction tuning data are often derived from datasets for downstream tasks, such as exam question answering Welbl et al. (2017), and further filtered or augmented by humans or existing LLMs (e.g., GPT-4 Achiam et al. (2023)).
General scientific LLMs are usually evaluated on common NLP tasks, such as named entity recognition (NER), relation extraction (RE) Luan et al. (2018), question answering (QA) Wang et al. (2024e), and classification Cohan et al. (2019).
2.2 Language + Graph
Beyond plain text, scientific papers are associated with rich metadata including venues, authors, and references Zhang et al. (2023g). Such metadata connect papers into a graph that complements text signals for characterizing paper semantics. To exploit metadata, some studies (Type 1.b, e.g., OAG-BERT Liu et al. (2022b)) concatenate paper text with venues/authors as input and perform MLM on both text and metadata; others (Type 3.a, e.g., SPECTER Cohan et al. (2020)) take citation links as supervision and train LLMs to encode linked papers closer in the embedding space. Recent approaches further modify the Transformer architecture in LLMs with Adapters Singh et al. (2023), GNN-nested Transformers Jin et al. (2023b), and Mixture-of-Experts Transformers Zhang et al. (2023f) to better capture graph signals.
Graph-aware scientific LLMs are often evaluated on tasks regarding the relation between two text units (e.g., paper-paper or query-paper), including link prediction, retrieval, recommendation, and author name disambiguation. SciDocs Cohan et al. (2020) and SciRepEval Singh et al. (2023) are widely adopted benchmark datasets.
2.3 Applications in Scientific Discovery
Performant scientific LLMs can work alongside researchers throughout the entire scientific discovery process. Leaving field-specific applications for later sections, here we underscore LLMs’ general usefulness in brainstorming and evaluation: Lahav et al. (2022) integrate LLMs into a search engine for the discovery of scientific challenges and directions; Wang et al. (2023e), Yang et al. (2024d), Baek et al. (2024), and Gu and Krenn (2024) leverage LLMs to generate novel scientific ideas, directions, and hypotheses on the basis of prior literature and existing knowledge; Zhang et al. (2023h) rely on LLMs to find expert reviewers for each submission; Liu and Shah (2023), Liang et al. (2024c), and D’Arcy et al. (2024) explore the capacity of GPT-4 to provide useful feedback on research papers to facilitate automatic review generation; Liang et al. (2024b, a) also observe the increasing use of LLMs in writing scientific papers and conference peer reviews.
3 LLMs in Mathematics (Table A2)
3.1 Language
The pre-training text corpora for math LLMs can be categorized into two classes: (1) multiple-choice QA, the representative datasets of which include MathQA Amini et al. (2019), Ape210K Zhao et al. (2020), and Math23K Wang et al. (2017); as well as (2) generative QA, the representative datasets of which include GSM8K Cobbe et al. (2021a), MATH Hendrycks et al. (2021), and MetaMathQA Yu et al. (2024c).
Similarly to general science LLMs, the backbone model of pioneering math LLMs is BERT (Type 1.a, e.g., GenBERT Geva et al. (2020) and MathBERT Shen et al. (2021)). These models are mostly trained via MLM, with the only exception being BERT-TD Li et al. (2022c), where a contrastive loss is adopted. For GPT-based math LLMs (Type 2.a, e.g., GSM8K-GPT Cobbe et al. (2021b) and NaturalProver Welleck et al. (2022)), we find a diversity of pre-training tasks: supervised fine-tuning, next token prediction, and instruction tuning. The most recent math LLMs (Type 2.a, e.g., Rho-Math Lin et al. (2024b) and MAmmoTH2 Yue et al. (2024c)) are based on LLaMA and are trained to follow natural language instructions. However, when an enormous pre-training dataset (e.g., 55 billion tokens) is available, next token prediction is still favored as the mere pre-training task Azerbayev et al. (2024); Lin et al. (2024b) or the companion task Shao et al. (2024); Ying et al. (2024) to build base models.
QA and math world problems (MWP) have been the most common evaluation tasks. In addition, quantitative reasoning contains more difficult problems, as the model has to provide a complete and self-contained solution without relying on external tools Shao et al. (2024); Lin et al. (2024b). We see a dominance of use from GSM8K and MATH for QA, and from MathQA and Math23K for MWP. For quantitative reasoning, MMLU-STEM Hendrycks et al. (2020) and Big-Bench Hard Suzgun et al. (2023) are the most widely adopted.
3.2 Language + Vision
Geometry is one of the most important branches of mathematics, and it expresses the settings jointly in text and diagrams. As such, it is mandatory to involve the vision modality for geometric LLMs. The most commonly used pre-training datasets for geometric LLMs include Geometry3K Lu et al. (2021) and GeoQA Chen et al. (2021), both of which contain multiple-choice geometry problems.
The key to incorporating the vision modality into LLMs is to encode the images and obtain linearized visual representations. Specifically, Inter-GPS Lu et al. (2021) (Type 2.d) uses RetinaNet Lin et al. (2017) to transform images into a set of relationships and then applies BART Lewis et al. (2020a) to produce the solution; G-LLaVA Gao et al. (2023) (Type 2.d) encodes visual input via a pre-trained vision Transformer (ViT), concatenates visual embeddings with textual embeddings, and then feeds the concatenation into LLaMA-2 Touvron et al. (2023b). These models are by default pre-trained via sequence-to-sequence tasks, where the problem is the input, and the ground-truth answer with optional rationale is the output. Auxiliary loss such as masked image modeling, image construction, or text-image matching, is optionally added for better visual modeling.
Geometric LLMs are evaluated through geometry problem solving, where the model is asked to select the correct answer given the diagram and its caption, the question, and answer options. Renowned evaluation datasets include Geometry3K Lu et al. (2021), GEOS Seo et al. (2015), and MathVista Lu et al. (2023b).
3.3 Table
A large proportion of math knowledge is stored in the form of tabular data. For the “Table” modality, notable resources for pre-training include WikiTableQuestions Pasupat and Liang (2015), WikiSQL Zhong et al. (2017), and WDC Web Table Lehmberg et al. (2016).
The challenge in tables is similar to that in diagrams, namely to obtain linearized table representations. In most cases, tables are squeezed into linear text sequences as part of the context and are prepended with the question text as the model input. As one of the first works in this line of research, TAPAS Herzig et al. (2020) (Type 1.a) adopts the MLM objective to predict the masked token in textual and tabular context. The most recent developments Li et al. (2024b); Zhang et al. (2024f) resemble the design of TableLlama Zhang et al. (2024e) (Type 2.b), with LLaMA-2 as the backbone and instruction tuning as the pre-training task.
Table LLMs are validated through table QA, where the model is asked to produce the correct answer given the table structure, data values, and a question text. Most existing studies have been evaluated on the WikiTableQuestions and WikiSQL datasets. TableInstruct Zhang et al. (2024e) is the most recently developed comprehensive benchmark integrating 14 datasets across 11 tasks.
3.4 Applications in Scientific Discovery
Math LLMs have great potential to assist humans in offering potential solutions. For instance, AlphaGeometry Trinh et al. (2024) combines an LLM with a symbolic deduction engine, where the LLM generates useful constructs and the symbolic engine applies formal logic to find solutions. AlphaGeometry solves 25 out of 30 classical geometry problems adapted from the International Mathematical Olympiad. Sinha et al. (2024) extend AlphaGeometry by adding Wu’s method Chou (1988), further solving 27 out of 30, surpassing human gold medalists. FunSearch Romera-Paredes et al. (2024) integrates LLM with program search. One notable achievement of FunSearch is its ability to find a new solution to the cap set problem in combinatorial optimization. The solutions generated can be faster and more efficient than those devised by human experts. In Li et al. (2024a), LLMs iteratively propose and critique statistical models by leveraging in-context learning and chain-of-thought reasoning Wei et al. (2022b).
4 LLMs in Physics (Table A3)
4.1 Language
As a derivative of BERT, astroBERT Grezes et al. (2021) (Type 1.a) is further pre-trained using astronomy-related papers via MLM and next sentence prediction. It is evaluated on the NER task. Likewise, AstroLLaMA Nguyen et al. (2023b) (Type 2.a) fine-tunes LLaMA-2 using over 300,000 astronomy abstracts from arXiv. It is evaluated on paper generation and recommendation tasks. AstroLLaMA-chat Perkowski et al. (2024) (Type 2.a) is the chat version of AstroLLaMA. It is continually trained on a GPT-4 generated domain-specific dialogue dataset. PhysBERT Hellert et al. (2024) (Type 1.a) is the first physics-specific model for sentence embedding trained on a curated corpus of physics literature based on 1.2 million physics papers on arXiv. It is evaluated on physics-tailored tasks, such as information retrieval, classification, and semantic similarity estimation.
4.2 Applications in Scientific Discovery
Transformer-based physics LLMs can potentially assist humans in solving differential equations and designing experiments. For instance, Cai et al. (2024) apply Transformer to predict the integer coefficients in the scattering amplitudes of Planar $\mathcal{N}=4$ Super Yang-Mills theory; RydbergGPT Fitzek et al. (2024) uses Transformer to learn the distribution of qubit measurement outcomes that describe an array of interacting Rydberg atoms; Arlt et al. (2024) present an initial trial that applies a code-generating LLM to synthesize experimental blueprints for a whole class of quantum systems in the form of Python code.
5 LLMs in Chemistry and Materials Science (Table A4)
5.1 Language
LLM pre-training corpora in chemistry and materials science typically come from research papers and databases (e.g., Materials Project Jain et al. (2013)). Besides, recent works adopt domain-specific instruction tuning datasets (e.g., Mol-Instructions Fang et al. (2023) and SMolInstruct Yu et al. (2024a)) derived from PubChem Kim et al. (2019), MoleculeNet Wu et al. (2018), etc.
Early studies on chemistry LLMs mostly adopt a moderate-sized encoder-only architecture pre-trained with MLM (Type 1.a, e.g., ChemBERT Guo et al. (2022), MatSciBERT Gupta et al. (2022), and BatteryBERT Huang and Cole (2022)). These models are usually evaluated on downstream tasks including reaction role labeling Guo et al. (2022) and abstract classification Gupta et al. (2022). Recently, researchers have focused more on large-scale decoder-only LLMs trained with next token prediction and instruction tuning (Type 2.a). Examples include ChemDFM Zhao et al. (2024), ChemLLM Zhang et al. (2024b), and LlaSMol Yu et al. (2024a). Given the desired generalization capability of such models, they are evaluated on a diverse set of tasks such as name conversion Kim et al. (2019), reaction prediction Jin et al. (2017), retrosynthesis Schneider et al. (2016), text-based molecule design Edwards et al. (2022), and crystal generation Antunes et al. (2023); Flam-Shepherd and Aspuru-Guzik (2023); Gruver et al. (2024).
5.2 Language + Graph
Graphs are appropriate data structures for characterizing molecules Jin et al. (2023a). Popular datasets containing molecular graphs include ChEBI-20 Edwards et al. (2021, 2022), ZINC Sterling and Irwin (2015), and PCDes Zeng et al. (2022).
In some scenarios, molecular graphs appear simultaneously with text information, thus existing works have explored how to encode both effectively. The first type of such models adopts a GNN as the graph encoder and an LLM as the text encoder. The two modalities are connected through contrastive learning Liu et al. (2023d) (Type 3.c). For example, Text2Mol Edwards et al. (2021) uses GCN Kipf and Welling (2016) and SciBERT to encode a molecule and its corresponding natural language description, respectively, for text-to-molecule retrieval. The second type of models utilizes an LLM to encode text and graphs simultaneously Zeng et al. (2022). Graphs can be either linearized to SMILES strings Edwards et al. (2022) (Type 2.c) or projected onto virtual tokens with graph encoders Zhao et al. (2023a); Liu et al. (2023e) (Type 2.d). For instance, 3D-MoLM Li et al. (2024c) uses a 3D molecular encoder to represent molecules as tokens, and feeds them together with instructions into LLaMA-2 for molecule-to-text retrieval and molecule captioning.
5.3 Language + Vision
Complementing text and graph modalities, molecular images form the vision modality in chemistry. Existing works adopt a similar philosophy to BLIP-2 Li et al. (2023b), which represents each image as tokens and feeds them into an LLM (Type 2.d). For example, GIT-Mol Liu et al. (2024a) projects all modalities, including graphs and images, into the latent text space and conducts encoding and decoding using T5 Raffel et al. (2020).
5.4 Molecule
Different from subsection 5.2, this subsection introduces models dealing with molecules without associated text information. That being said, comparable approaches inspired by LLMs are utilized to develop molecular language models Flam-Shepherd et al. (2022). To be specific, most studies adopt SMILES or SELFIES Krenn et al. (2020) strings as the sequential representation of molecules. Similar to the trend in the “Language” modality, pioneering molecular LLMs focus on representation learning with bidirectional Transformer encoders (Type 1.c, e.g., SMILES-BERT Wang et al. (2019) and MoLFormer Ross et al. (2022)). For instance, ChemBERTa Chithrananda et al. (2020) adopts the architecture and pre-training strategy similar to those of RoBERTa Liu et al. (2019). These models exhibit extraordinary abilities in molecular understanding tasks such as molecular property prediction (e.g., toxicity classification Wu et al. (2018) and atomization energy regression Ramakrishnan et al. (2014)) as well as virtual screening Riniker and Landrum (2013). Later works explore representing molecules in an autoregressive fashion (Type 2.c, e.g., BARTSmiles Chilingaryan et al. (2024) and ChemGPT Frey et al. (2023)). For instance, T5Chem Lu and Zhang (2022) adopts the T5 backbone and a sequence-to-sequence pre-training objective. These models are evaluated in generative tasks that include molecule generation Gaulton et al. (2017), reaction prediction, and retrosynthesis. Besides linearizing molecules, there are studies modifying the Transformer architecture to admit molecular graphs, such as MAT Maziarka et al. (2020) and R-MAT Maziarka et al. (2024).
5.5 Applications in Scientific Discovery
Previous studies have shown that LLMs facilitate autonomous chemical research. For example, Bran et al. (2024) present a chemistry LLM agent, ChemCrow, that can integrate expert-designed tools for organic synthesis, drug discovery, and materials design; Zheng et al. (2023a) demonstrate that LLMs can perform knowledge synthesis from the scientific literature, knowledge inference from data, and interpretable explanation generation in chemistry; Boiko et al. (2023) develop an LLM-empowered intelligence system, Coscientist, that can design, plan, and perform chemical research. Moreover, LLMs accomplish complex tasks in chemistry, such as drug and catalyst design and molecular discovery, purely from instructions White (2023). For instance, Ramos et al. (2023) study catalyst and molecule design with in-context learning, removing the requirement for traditional training or simulation processes; ChatDrug Liu et al. (2024b) explores drug editing using LLMs with a prompt module, a domain feedback module, and a conversation module; Jablonka et al. (2024) find that fine-tuned LLMs perform comparably to, or even better than, conventional techniques for many chemistry applications, spanning from the properties of molecules and materials to the yield of chemical reactions; DrugAssist Ye et al. (2023a) serves as an LLM-based interactive model for molecule optimization through human-machine dialogue; Sprueill et al. (2023, 2024) use LLMs as agents to search for effective catalysts through Monte Carlo Tree Search and the feedback from an atomistic neural network model; Wang et al. (2024b) re-engineer crossover and mutation operations for molecular discovery using LLMs trained on extensive chemical datasets. Meanwhile, benchmarking studies by Mirza et al. (2024) demonstrate that although LLMs achieve superhuman proficiency in many chemical tasks, further research is critical to enhancing their safety and utility in chemical sciences.
6 LLMs in Biology and Medicine (Table A5)
6.1 Language
Besides research articles (e.g., titles/abstracts from PubMed Lu (2011) and full text from PMC Beck and Sequeira (2003)), pre-training corpora for biomedical LLMs include electronic health records (e.g., MIMIC-III Johnson et al. (2016), MIMIC-IV Johnson et al. (2023)), knowledge bases (e.g., UMLS Bodenreider (2004)), and health-related social media posts (e.g., COVID-19 tweets Müller et al. (2023)). Recent studies further collect supervised fine-tuning and preference optimization datasets from medical exam questions, knowledge graphs, and doctor-patient dialogues. Examples include ChiMed Ye et al. (2023b), MedInstruct-52k Zhang et al. (2023d), and BiMed1.3M Acikgoz et al. (2024), many of which have non-English components (e.g., Chinese and Arabic).
The watershed moment in the evolution biomedical LLMs is still the emergence of billion-parameter architectures and instruction tuning. Before that, a wide variety of moderate-sized backbones are explored, including both encoder-based (Type 1.a, e.g., BioBERT Lee et al. (2020), Bio-ELECTRA Ozyurt (2020), BioRoBERTa Lewis et al. (2020b), BioALBERT Naseem et al. (2022), and Clinical-Longformer Li et al. (2022a)) and (encoder-) decoder-based ones (Type 2.a, e.g., SciFive Phan et al. (2021), BioBART Yuan et al. (2022a), and BioGPT Luo et al. (2022)). Evaluation tasks for these models range from biomedical NER, RE, sentence similarity estimation, document classification, and QA (i.e., the BLURB benchmark Gu et al. (2021)) to natural language inference (NLI) Romanov and Shivade (2018) and entity linking Doğan et al. (2014). After the watershed, the trend becomes instruction-tuning billion-parameter LLMs (Type 2.a, e.g., Med-PaLM Singhal et al. (2023a), MedAlpaca Han et al. (2023), and BioMistral Labrak et al. (2024)). Accordingly, evaluation tasks now include single-round QA Jin et al. (2021); Pal et al. (2022) and multi-round dialogue Wang et al. (2023f). Meanwhile, there are studies proposing a Bi-Encoder architecture (Type 3.a, e.g., Jin et al. (2023c) and Xu et al. (2024)) that specifically targets biomedical retrieval tasks, the benchmarks of which are NFCorpus Boteva et al. (2016), TREC-COVID Voorhees et al. (2021), etc.
6.2 Language + Graph
Biomedical ontologies capture rich types of relations between entities. Analogously, citation links characterize connections between biomedical papers. Intuitively, jointly leveraging text and such graph information paves the way for multi-hop reasoning in QA. For instance, Yasunaga et al. (2022a) propose to use an LLM and a GNN to encode text and ontology signals, respectively, and deeply fuse them (Type 3.c); Yasunaga et al. (2022b) concatenate text segments from two linked papers together and feed the sequence into an LLM for pre-training, which is essentially appending a metadata neighbor (i.e., reference) as context for MLM (Type 1.b). Both approaches demonstrate significant improvement in QA tasks that require complex reasoning.
6.3 Language + Vision
Biomedical text-image pairs typically come from two sources: (1) medical reports, such as chest X-rays (e.g., MIMIC-CXR Johnson et al. (2019)) and pathology reports Huang et al. (2023b); as well as (2) figure-caption pairs extracted from biomedical papers (e.g., ROCO Pelka et al. (2018) and MedICaT Subramanian et al. (2020)).
Most biomedical vision-language models exploit the CLIP architecture Radford et al. (2021), where a text encoder and an image encoder are jointly trained to map the paired text and image closer via contrastive learning (Type 3.d). The choice of the text encoder evolves from BERT Zhang et al. (2022) and GPT-2 Huang et al. (2023b) to LLaMA Wu et al. (2023) and LLaMA-2 Liu et al. (2023b), while the image encoder evolves from ResNet Huang et al. (2021) to ViT Zhang et al. (2023c) and Swin Transformer Thawkar et al. (2024). MLM, masked image modeling, and text-text/image-image contrastive learning (i.e., by creating augmented views within the language/vision modality) are sometimes adopted as auxiliary pre-training tasks. Besides CLIP, other general-domain vision-language architectures, such as LLaVA Li et al. (2023a), PaLM-E Tu et al. (2024), and Gemini Saab et al. (2024), have been explored. For instance, LLaVA-Med (Type 2.d) encodes images onto several visual tokens and prepends them to text tokens as the LLM input. Evaluation tasks of these models encompass image classification, segmentation, object detection, vision QA, text-to-image/image-to-text retrieval, and report generation, the benchmarks of which include CheXpert Irvin et al. (2019), PadChest Bustos et al. (2020), SLAKE Liu et al. (2021a), etc.
6.4 Protein, DNA, RNA, and Multiomics
The FASTA format Lipman and Pearson (1985) naturally represents proteins as amino acid sequences and DNAs/RNAs as nucleotide sequences, enabling models to treat them as “languages”. Representative resources of such sequences include UniRef Suzek et al. (2015) and Swiss-Prot Bairoch and Apweiler (2000) for proteins, GRCh38 Harrow et al. (2012) and the 1000 Genomes Project Consortium (2015) for DNAs, as well as RNAcentral Consortium (2019) for RNAs.
Encoder-only protein, DNA, and RNA LLMs (Type 1.d), such as ESM-2 Lin et al. (2023b), DNABERT Ji et al. (2021), and RNABERT Akiyama and Sakakibara (2022), adopt BERT-like architectures and MLM as the pre-training task (i.e., predicting masked amino acids, nucleotides, $k$-mers, or codons); decoder-only models, such as ProGen Madani et al. (2023) and DNAGPT Zhang et al. (2023a), exploit GPT-like architectures and next token prediction as the pre-training task. There are also studies jointly considering text and protein modalities. For instance, ProtST Xu et al. (2023b) matches protein sequences with their text descriptions (i.e., names and functions) via contrastive learning (Type 3.b); BioMedGPT Luo et al. (2023c) first projects proteins onto tokens and then inputs these tokens together with text into LLaMA-2 for instruction tuning, bearing similarity with Type 2.d.
Existing multiomics LLMs mainly focus on single-cell transcriptomics (e.g., scRNA-seq) data, such as the expression levels of genes within a single cell Franzén et al. (2019). Besides BERT-based (e.g., Geneformer Theodoris et al. (2023)) and GPT-based (e.g., scGPT Cui et al. (2024)) architectures, Performer Yang et al. (2022a); Hao et al. (2024) is widely used due to its linear attention complexity in handling long scRNA-seq data.
6.5 Applications in Scientific Discovery
Similarly to chemistry, LLMs can automate experiments in biological and medical research. For example, CRISPR-GPT Huang et al. (2024) augments an LLM agent with domain knowledge to enhance the design process of CRISPR-based gene-editing experiments; TrialMind Wang et al. (2024f) utilizes LLMs to extract results and synthesize clinical evidence from the literature for medical discovery. Moreover, LLMs can encode biological sequences to capture structural properties, guide protein design, and evaluate the evolutionary fitness of viral variants. For instance, ESM-1b Rives et al. (2021) and ESM-2 Lin et al. (2023b) enable accurate structure prediction of proteins without expensive and time-consuming experiments; Ferruz and Höcker (2022) fine-tune LLMs on protein families, which can generate highly divergent but still potentially functional novel sequences; Hie et al. (2021) develop LLMs that can predict viral escape mutations.
7 LLMs in Geography, Geology, and Environmental Science (Table A6)
7.1 Language
Geoscience research papers, climate-related news articles, Wikipedia pages, corporate sustainability reports, knowledge bases (e.g., GAKG Deng et al. (2021)), and point-of-interest (POI) data (e.g., OpenStreetMap Haklay and Weber (2008)) constitute the pre-training corpora for geoscience LLMs.
Preliminary research on geoscience LLMs focuses on pre-training bidirectional LLMs with the Transformer encoder backbone (Type 1.a, e.g., ClimateBERT Webersinke et al. (2021), SpaBERT Li et al. (2022b), and MGeo Ding et al. (2023)). For instance, SpaBERT and MGeo perform MLM on a sequence of geolocations for geographic entity linking and query-POI matching, respectively. More recently, related studies concentrate on scaling up decoding-style autoregressive LLMs in geoscience (Type 2.a, e.g., K2 Deng et al. (2024), OceanGPT Bi et al. (2023b), and GeoGalactica Lin et al. (2024c)). For instance, K2 and OceanGPT adapt LLaMA to geoscience and ocean science, respectively, via supervised fine-tuning with domain-specific instructions curated by human experts and/or augmented by general-domain LLMs. Evaluations of such models are conducted on geoscience benchmarks, such as GeoBench Deng et al. (2024) and OceanBench Bi et al. (2023b), which encompass a broad range of tasks including QA, classification, knowledge probing, reasoning, summarization, and generation.
7.2 Language + Graph
Some geoscience applications involve graph signals, such as heterogeneous POI networks and knowledge graphs. To handle such signals and text jointly, ERNIE-GeoL Huang et al. (2022) introduces a Transformer-based aggregation layer to deeply fuse text and POI information within a BERT-based architecture; PK-Chat Deng et al. (2023) combines an LLM with a pointer generation network on a knowledge graph to build a knowledge-driven dialogue system.
7.3 Language + Vision
Aerial views, together with location descriptions, profile urban regions. To address language and vision modalities jointly, UrbanCLIP Yan et al. (2024) considers the CLIP architecture (Type 3.d), which is also widely adopted by biomedical vision-language models as mentioned in subsection 6.3, to perform text-image contrastive learning for urban indicator prediction.
7.4 Climate Time Series
The intuitions and methodologies used in LLMs also facilitate the construction of climate foundation models. Based on the ERA5 Hersbach et al. (2020) and CMIP6 Eyring et al. (2016) datasets of climate time series, previous studies exploit the ViT and Swin Transformer architectures to pre-train foundation models for weather forecasting. Representative models include FourCastNet Pathak et al. (2022), Pangu-Weather Bi et al. (2023a), etc.
7.5 Applications in Scientific Discovery
In geography, Wang et al. (2023b) and Zhou et al. (2024b) highlight the potential of LLMs in urban planning from sustainability, living, economic, disaster, and environmental perspectives. In geology, besides climate and weather forecasting, foundation models have been applied to simultaneous earthquake detection and phase picking Mousavi et al. (2020). In environmental science, ChatClimate Vaghefi et al. (2023) enhances GPT-4 by providing access to external, scientifically accurate knowledge on climate change to build a climate science conversational AI.
8 Challenges and Future Directions
In this survey, we compile literature that elucidates the data, architectures, and tasks used for scientific LLM pre-training, as well as how scientific LLMs have been applied to downstream applications in scientific discovery. In particular, we underscore analogous architectures, tasks, and trends observed during the evolution of scientific LLMs across different fields and modalities. Beyond reviewing prior research, we present several challenges to inspire further exploration of this topic.
Diving into Fine-Grained Themes. Most existing scientific LLMs target a coarse-grained field (e.g., chemistry), while some tasks rely on highly specialized knowledge of a fine-grained theme (e.g., Suzuki coupling). When LLMs are pre-trained on more general corpora, frequently appeared signals may dominate the model parameter space, and domain-specific tail knowledge may be wiped out. We believe that automatically curating in-depth, theme-focused knowledge graphs Hope et al. (2021) to guide the generation process will be a promising direction to tackle this issue.
Generalizing to Out-of-Distribution Scientific Data. In the scientific domain, it is common that the testing distribution shifts from the training distribution Zhang et al. (2023e): novel scientific concepts keep emerging in newly published papers; unseen molecules with different scaffolds and unseen proteins with different numbers of peptide chains may appear during testing. Handling such out-of-distribution data remains a challenge for pre-trained scientific LLMs. To our knowledge, invariant learning Arjovsky et al. (2019) can serve as the theoretical foundation for out-of-distribution analyses, and how to integrate it into LLM pre-training is worth exploring.
Facilitating Trustworthy Predictions. LLMs can generate plausible-sounding but factually incorrect output, commonly known as hallucination Ji et al. (2023), which is particularly dangerous in high-stakes scientific domains such as chemistry and biomedicine. To mitigate this issue, retrieval-augmented generation (RAG) provides LLMs with relevant, up-to-date, and trustworthy information. However, previous RAG studies in the scientific domain mainly focus on retrieving text Xiong et al. (2024) and knowledge Jin et al. (2024), while scientific data are heterogeneous and multi-modal. We envision that cross-modal RAG (e.g., guiding text generation with relevant chemicals and proteins) will present additional opportunities to further enhance the trustworthiness of scientific LLMs.
Limitations
This survey primarily covers LLMs in mathematics and natural sciences. We are aware that LLMs can also significantly impact social sciences by achieving remarkable performance in representative tasks Ziems et al. (2024) and serving as agents for social simulation experiments Horton (2023), but we leave the survey of these efforts as future work due to space limitations. In addition, this paper focuses on LLMs pre-trained on scientific data or augmented with domain-specific knowledge to benefit scientific discovery. There are studies Guo et al. (2023); Wang et al. (2024e); Yue et al. (2024a); Liang et al. (2024d) proposing new benchmark datasets of scientific problems but evaluating the performance of general-purpose LLMs only, and we do not include these works in our survey. Furthermore, some LLMs may belong to more than one field or modality category given our classification criteria in the paper. For instance, BioMedGPT Luo et al. (2023c) is pre-trained on biology and chemistry data jointly; GIT-Mol Liu et al. (2024a) considers the language, graph, and vision modalities simultaneously. For the sake of brevity, we introduce each of them in only one subsection.
Acknowledgments
Research was supported in part by US DARPA KAIROS Program No. FA8750-19-2-1004 and INCAS Program No. HR001121C0165, National Science Foundation IIS-19-56151, and the Molecule Maker Lab Institute: An AI Research Institutes program supported by NSF under Award No. 2019897, and the Institute for Geospatial Understanding through an Integrative Discovery Environment (I-GUIDE) by NSF under Award No. 2118329. Any opinions, findings, and conclusions or recommendations expressed herein are those of the authors and do not necessarily represent the views, either expressed or implied, of DARPA or the U.S. Government.
Appendix A Summary Tables of Scientific LLMs
Table A1-Table A6 summarize the modality, number of parameters, model architecture, pre-training data, pre-training task(s), and evaluation task(s) of scientific LLMs in each field. Within each field, we categorize models according to their modality; within each modality, we sort models chronologically. To be specific, if a paper has a preprint (e.g., arXiv or bioRxiv) version, its publication date is according to the preprint service. Otherwise, its publication date is according to the conference proceeding or journal.
Table A1: Summary of LLMs in general science. “L”: Language; “L+G”: Language + Graph; $\sim$”: generally adopting the architecture but with modifications; “MLM”: masked language modeling; “NSP”: next sentence prediction; “NER”: named entity recognition; “RE”: relation extraction; “QA”: question answering.
Model Modality Size Architecture Pre-training Data Pre-training Task(s) Evaluation Task(s)
SciBERT Beltagy et al. (2019) L 110M BERT Semantic Scholar MLM, NSP NER, RE, classification, parsing
SciGPT2 Luu et al. (2021) L 117M GPT-2 S2ORC next token prediction paper relationship explanation
CATTS Cachola et al. (2020) L 406M BART SciTLDR sequence to sequence summarization
SciNewsBERT Smeros et al. (2021) L 110M BERT news headlines MLM, NSP scientific claim extraction
ScholarBERT Hong et al. (2023) L 340M, 770M BERT Public.Resource.Org, MLM NER, RE, classification
Wikipedia, BookCorpus
AcademicRoBERTa Yamauchi et al. (2022) L 125M RoBERTa CiNii MLM classification,
author identification
Galactica Taylor et al. (2022) L 125M, 1.3B, Galactica papers, code, next token prediction, QA, link prediction,
6.7B, 30B, reference materials, instruction tuning knowledge probing,
120B knowledge bases, quantitative reasoning,
web crawl data, chemical name conversion,
instructions molecule classification,
protein function prediction
DARWIN Xie et al. (2023) L 7B LLaMA papers, QA pairs, instruction tuning QA, classification, regression
instructions
FORGE Yin et al. (2023) L 1.4B, 13B, GPT-NeoX CORE, AMiner, MAG, next token prediction QA, classification, regression
22B SCOPUS, arXiv
SciGLM Zhang et al. (2024a) L 6B, 32B ChatGLM SciInstruct instruction tuning QA, quantitative reasoning
SPECTER Cohan et al. (2020) L+G 110M BERT Semantic Scholar link prediction classification, link prediction,
recommendation
[…]
OAG recommendation, retrieval,
author name disambiguation
ASPIRE Mysore et al. (2022) L+G 110M BERT S2ORC link prediction paper similarity estimation
SciNCL Ostendorff et al. (2022) L+G 110M BERT Semantic Scholar link prediction classification, link prediction,
recommendation
SPECTER 2.0 Singh et al. (2023) L+G 113M Adapters SciRepEval classification, regression, classification, regression,
link prediction, retrieval link prediction, retrieval,
author name disambiguation,
paper-reviewer matching
SciPatton Jin et al. (2023b) L+G GraphFormers MAG MLM, link prediction classification, link prediction
SciMult Zhang et al. (2023f) L+G 138M MoE MAG, classification, classification, link prediction,
Semantic Scholar, link prediction, retrieval recommendation, retrieval,
SciRepEval patient-article/patient matching
Table A2: Summary of LLMs in mathematics. “L+V”: Language + Vision; “MWP”: math word problems. Other notations have the same meaning as in previous tables.
Model Modality Size Architecture Pre-training Data Pre-training Task(s) Evaluation Task(s)
GenBERT Geva et al. (2020) L 110M BERT Wikipedia MLM, QA, MWP
sequence to sequence
MathBERT Shen et al. (2021) L 110M BERT arXiv, math curricula, MLM classification, auto-grading
syllabi, textbooks
MWP-BERT Liang et al. (2022) L 110M BERT Ape210K MLM, regression, QA, MWP
classification
BERT-TD Li et al. (2022c) L 110M BERT Math23K, MathQA sequence to sequence, QA, MWP
contrastive learning
GSM8K-GPT Cobbe et al. (2021b) L 6B, 175B GPT-3 GSM8K supervised fine-tuning QA, MWP
DeductReasoner Jie et al. (2022) L 125M RoBERTa MAWPS, Math23K, sequence to sequence QA, MWP
MathQA, SVAMP
NaturalProver Welleck et al. (2022) L 175B GPT-3 NaturalProofs supervised fine-tuning mathematical proof generation
Minerva Lewkowycz et al. (2022) L 8B, 62B, PaLM arXiv, math web pages next token prediction QA, MWP, quantitative reasoning
540B
[…]
WizardMath Luo et al. (2023a) L 7B, 13B, LLaMA-2 GSM8K, MATH instruction tuning QA, MWP
70B
MAmmoTH Yue et al. (2024b) L 7B, 13B, LLaMA-2 MathInstruct instruction tuning QA, MWP
34B, 70B
7B Mistral
MetaMath Yu et al. (2024c) L 7B, 13B, LLaMA-2 MetaMathQA instruction tuning QA, MWP
70B
7B Mistral
ToRA Gou et al. (2024) L 7B, 13B, LLaMA-2 ToRA-Corpus instruction tuning QA, MWP
34B, 70B
MathCoder Wang et al. (2024c) L 7B, 13B, LLaMA-2 MathCodeInstruct instruction tuning QA, MWP
34B, 70B
Llemma Azerbayev et al. (2024) L 7B, 34B LLaMA-2 Proof-Pile-2 next token prediction QA, MWP, quantitative reasoning
OVM Yu et al. (2024b) L 7B LLaMA-2 GSM8K supervised fine-tuning QA, MWP, quantitative reasoning
7B Mistral
DeepSeekMath Shao et al. (2024) L 7B DeepSeek math web pages, next token prediction, QA, MWP, quantitative reasoning,
instructions instruction tuning formal translation
InternLM-Math Ying et al. (2024) L 7B, 20B InternLM2 Knowledge Pile, next token prediction, QA, MWP, quantitative reasoning,
Proof-Pile-2, instruction tuning formal translation
instructions
OpenMath Toshniwal et al. (2024) L 7B, 13B, LLaMA-2 OpenMathInstruct-1 instruction tuning QA, MWP
34B, 70B
7B Mistral
[…]
7B Mistral SlimPajama,
StarCoderData
MAmmoTH2 Yue et al. (2024c) L 8B LLaMA-3 WebInstruct instruction tuning QA, MWP, quantitative reasoning
7B Mistral
8$\times$7B Mixtral
[…]
RetinaNet
Geoformer Chen et al. (2022a) L+V VL-T5 + UniGeo sequence to sequence geometry problem solving
ResNet
SCA-GPS Ning et al. (2023) L+V RoBERTa + GeoQA, Geometry3K masked image modeling, geometry problem solving
ViT sequence to sequence
(Mathematics, Table Continued)
Model Modality Size Architecture Pre-training Data Pre-training Task(s) Evaluation Task(s)
UniMath-Flan-T5 L+V Flan-T5 + SVAMP, GeoQA, image reconstruction, MWP,
     Liang et al. (2023) VQ-VAE TabMWP sequence to sequence geometry problem solving
G-LLaVA Gao et al. (2023) L+V 7B, 13B LLaVA GeoQA+, Geometry3K text-image matching, geometry problem solving
instruction tuning
TAPAS Herzig et al. (2020) Table 110M, 340M BERT Wikipedia MLM table QA
TaBERT Yin et al. (2020) Table 110M, 340M BERT Wikipedia, MLM, table QA
WDC Web Table cell value recovery
GraPPa Yu et al. (2021) Table 355M RoBERTa Wikipedia MLM, table QA
SQL semantic prediction
TUTA Wang et al. (2021) Table 110M BERT Wikipedia, MLM, cell type classification,
WDC Web Table, cell-level cloze, table type classification
spreadsheets table context retrieval
RCI Glass et al. (2021) Table 12M ALBERT WikiSQL, TabMCQ, classification table QA
WikiTableQuestions
TABBIE Iida et al. (2021) Table 110M ELECTRA Wikipedia, VizNet MLM, column/row population,
replaced token detection column type classification
TAPEX Liu et al. (2022a) Table 140M, 406M BART WikiTableQuestions sequence to sequence table QA
FORTAP Cheng et al. (2022) Table 110M BERT spreadsheets MLM, table QA,
numerical reference prediction, formula prediction,
numerical calculation prediction cell type classification
OmniTab Jiang et al. (2022) Table 406M BART Wikipedia sequence to sequence table QA
ReasTAP Zhao et al. (2022) Table 406M BART Wikipedia sequence to sequence table QA, table fact verification,
table-to-text generation
Table-GPT Li et al. (2024b) Table 175B GPT-3.5 instructions instruction tuning table QA, column-finding,
ChatGPT missing-value identification,
column type classification,
data transformation,
table matching, data cleaning
TableLlama Zhang et al. (2024e) Table 7B LLaMA-2 TableInstruct instruction tuning table QA, RE, entity linking,
column type classification,
column/row population,
table fact verification,
cell description
TableLLM Zhang et al. (2024f) Table 7B, 13B LLaMA-2 WikiTQ, FeTaQA, instruction tuning table QA, table updating,
TAT-QA, WikiSQL, table merging, table charting
Spider
Table A3: Summary of LLMs in physics. Notations have the same meaning as in previous tables.
Model Modality Size Architecture Pre-training Data Pre-training Task(s) Evaluation Task(s)
astroBERT Grezes et al. (2021) L 110M BERT NASA Astrophysics Data System MLM, NSP NER
AstroLLaMA Nguyen et al. (2023b) L 7B LLaMA-2 arXiv next token prediction paper generation,
paper similarity estimation
AstroLLaMA-Chat Perkowski et al. (2024) L 7B LLaMA-2 QA pairs, LIMA, OpenOrca, UltraChat instruction tuning QA
PhysBERT Hellert et al. (2024) L 110M BERT arXiv MLM, classification, retrieval,
contrastive learning clustering
Table A4: Summary of LLMs in chemistry and materials science. “L+G+V”: Language + Graph + Vision; “KG”: knowledge graph; “SMILES”: simplified molecular-input line-entry system. Other notations have the same meaning as in previous tables.
Model Modality Size Architecture Pre-training Data Pre-training Task(s) Evaluation Task(s)
ChemBERT Guo et al. (2022) L 110M BERT chemistry journals MLM NER
MatSciBERT Gupta et al. (2022) L 110M BERT ScienceDirect MLM NER, RE, classification
MatBERT Trewartha et al. (2022) L 110M BERT materials science journals MLM NER
BatteryBERT Huang and Cole (2022) L 110M BERT Elsevier, Springer, RSC MLM QA, classification
MaterialsBERT Shetty et al. (2023) L 110M BERT materials science journals MLM, NSP NER
Recycle-BERT Kumar et al. (2023) L 110M BERT plastic recycling articles classification QA, classification
CatBERTa Ock et al. (2023) L 125M RoBERTa OC20 regression regression
LLM-Prop Rubungo et al. (2023) L 37M T5 (encoder) Materials Project classification, regression classification, regression
ChemDFM Zhao et al. (2024) L 13B LLaMA chemistry papers, next token prediction, QA, classification,
textbooks, instructions instruction tuning name conversion,
molecule captioning,
text-based molecule design,
reaction prediction, retrosynthesis
CrystalLLM Gruver et al. (2024) L 7B, 13B, LLaMA-2 Materials Project instruction tuning crystal generation
70B
ChemLLM Zhang et al. (2024b) L 7B InternLM2 QA pairs, ChemData instruction tuning QA, classification,
name conversion,
molecule captioning,
text-based molecule design,
reaction prediction, retrosynthesis
LlaSMol Yu et al. (2024a) L 6.7B Galactica SMolInstruct instruction tuning QA, classification, regression,
7B LLaMA-2 name conversion,
7B Mistral molecule captioning,
text-based molecule design,
reaction prediction, retrosynthesis
Text2Mol Edwards et al. (2021) L+G BERT + PubChem, ChEBI-20 text-graph matching text-to-molecule retrieval
GCN
KV-PLM Zeng et al. (2022) L+G 110M BERT S2ORC, PubChem text-graph matching NER, RE, classification,
text-to-molecule retrieval,
molecule-to-text retrieval
MolT5 Edwards et al. (2022) L+G 60M, 220M, T5 C4, ZINC, ChEBI-20 sequence to sequence molecule captioning,
770M text-based molecule design
MoMu Su et al. (2022) L+G BERT + S2ORC, PubChem text-graph matching classification,
GIN text-to-molecule retrieval,
molecule-to-text retrieval,
molecule captioning,
text-based molecule design
(Chemistry and Materials Science, Table Continued)
Model Modality Size Architecture Pre-training Data Pre-training Task(s) Evaluation Task(s)
MoleculeSTM Liu et al. (2023d) L+G BERT + PubChem text-graph matching classification,
GIN text-to-molecule retrieval,
molecule-to-text retrieval,
text-based molecule design
Text+Chem T5 L+G 60M, 220M T5 Pistachio, ChEBI-20, sequence to sequence molecule captioning,
     Christofidellis et al. (2023) experimental procedures text-based molecule design,
reaction prediction, retrosynthesis,
paragraph-to-action generation
[…]
GIN text-graph matching text-to-molecule retrieval,
molecule-to-text retrieval,
molecule captioning,
text-based molecule design
MolCA Liu et al. (2023e) L+G Galactica + PubChem text-graph matching, classification, name conversion,
GIN graph-to-text generation molecule-to-text retrieval,
molecule captioning,
functional group counting
InstructMol Cao et al. (2023) L+G LLaMA + PubChem, MoleculeNet, text-graph matching, classification, regression,
GIN ChEBI-20, USPTO instruction tuning molecule captioning,
reaction prediction, retrosynthesis,
reagent selection
3D-MoLM Li et al. (2024c) L+G LLaMA-2 + PubChem, 3D-MoIT text-graph matching, QA, regression,
Uni-Mol graph-to-text generation, molecule-to-text retrieval,
instruction tuning molecule captioning
GIT-Mol Liu et al. (2024a) L+G+V BERT + PubChem, ChEBI-20 text-graph/image/text classification,
GIN + matching, molecule captioning,
Swin supervised fine-tuning text-based molecule design,
molecule image recognition
[…]
ChemBERTa Chithrananda et al. (2020) Molecule 125M RoBERTa PubChem MLM classification
MolBERT Fabian et al. (2020) Molecule 110M BERT ChEMBL MLM, regression, classification, regression,
SMILES equivalence virtual screening
rxnfp Schwaller et al. (2021b) Molecule 110M BERT Pistachio, USPTO classification classification,
reaction representation learning
[…]
MoLFormer Ross et al. (2022) Molecule 47M linear PubChem, ZINC MLM classification, regression
attention
[…]
MoleculeNet regression reaction prediction, retrosynthesis,
molecule generation
[…]
regression
[…]
reaction prediction, retrosynthesis
[…]
1.2B
Uni-Mol Zhou et al. (2023) Molecule SE(3) ZINC, ChEMBL, 3D position recovery classification, regression,
Transformer RCSB PDB molecule conformation generation,
binding pose prediction
[…]
polyBERT Molecule 86M DeBERTa density functional theory, MLM, regression
     Kuenneth and Ramprasad (2023) experiments regression
[…]
     Abdel-Aty and Gould (2022) PubChem, ChEMBL, virtual screening
USPTO
[…]
SMILES-property reaction prediction, retrosynthesis,
matching SMILES-to-property generation,
property-to-SMILES generation
BARTSmiles Chilingaryan et al. (2024) Molecule 406M BART ZINC sequence to sequence classification, regression,
reaction prediction, retrosynthesis
MolGen Fang et al. (2024) Molecule 406M BART ZINC, NPASS sequence to sequence, molecule generation
7B LLaMA prefix tuning
[…]
PolyNC Qiu et al. (2024a) Molecule 220M T5 density functional theory, sequence to sequence classification, regression
experiments
Table A5: Summary of LLMs in biology and medicine. “Multi”: Multiomics (e.g., single-cell); “NLI”: natural language inference; “VQA”: visual question answering; “EHR”: electronic health record; “EMR”: electronic medical record; “PPI”: protein-protein interaction. Other notations have the same meaning as in previous tables.
Model Modality Size Architecture Pre-training Data Pre-training Task(s) Evaluation Task(s)
BioBERT Lee et al. (2020) L 110M, 340M BERT PubMed, PMC MLM, NSP NER, RE, QA
BioELMo Jin et al. (2019) L 93M ELMo PubMed next token prediction, NER, NLI
previous token prediction
ClinicalBERT L 110M BERT MIMIC-III MLM, NSP NER, NLI
     Alsentzer et al. (2019)
ClinicalBERT Huang et al. (2019) L 110M BERT MIMIC-III next token prediction, word similarity estimation,
previous token prediction hospital readmission prediction
BlueBERT Peng et al. (2019) L 110M, 340M BERT PubMed, MIMIC-III MLM, NSP NER, RE, NLI, classification,
sentence similarity estimation
[…]
Research Datalink
[…]
Clinical XLNet Huang et al. (2020) L 110M XLNet MIMIC-III permutation language modeling mortality prediction
ouBioBERT Wada et al. (2020) L 110M BERT PubMed MLM, NSP NER, RE, NLI, classification,
sentence similarity estimation
COVID-Twitter-BERT L 340M BERT COVID-19 tweets MLM, NSP classification, sentiment analysis,
     Müller et al. (2023) stance prediction
[…]
(Biology and Medicine, Table Continued)
Model Modality Size Architecture Pre-training Data Pre-training Task(s) Evaluation Task(s)
Bio-ELECTRA Ozyurt (2020) L 110M ELECTRA PubMed MLM, replaced token detection NER, QA
BiomedBERT Gu et al. (2021) L 110M, 340M BERT PubMed, PMC MLM, NSP NER, RE, QA, classification,
sentence similarity estimation
MCBERT Zhang et al. (2020) L 110M BERT Chinese media, MLM, NSP NER, QA, classification, retrieval,
encyclopedia, EHRs paraphrase identification
[…]
BioRedditBERT L 110M BERT Reddit entity linking entity linking
     Basaldella et al. (2020)
BioMegatron Shin et al. (2020) L 345M BERT PubMed, PMC MLM, NSP NER, RE, QA
SapBERT Liu et al. (2021b) L 110M BERT UMLS synonym alignment entity linking
ClinicalTransformer L 110M BERT MIMIC-III MLM, NSP, NER
     Yang et al. (2020) 125M RoBERTa sentence order prediction,
12M ALBERT replaced token detection,
110M ELECTRA permutation language modeling
110M XLNet
149M Longformer
86M DeBERTa
BioRoBERTa Lewis et al. (2020b) L 125M, 355M RoBERTa PubMed, PMC, MLM NER, RE, NLI, classification
MIMIC-III
RAD-BERT Bressem et al. (2020) L 110M BERT radiology reports MLM, NSP classification
BioMedBERT L 340M BERT BREATHE MLM, NSP NER, RE, QA, retrieval
     Chakraborty et al. (2020)
[…]
ELECTRAMed Miolo et al. (2021) L 110M ELECTRA PubMed MLM, replaced token detection NER, RE, QA
KeBioLM Yuan et al. (2021) L 110M BERT PubMed, UMLS MLM, NER, entity linking NER, RE, knowledge probing
SciFive Phan et al. (2021) L 220M, 770M T5 PubMed, PMC sequence to sequence NER, RE, QA, NLI, classification
BioALBERT Naseem et al. (2022) L 12M, 18M ALBERT PubMed, PMC, MLM, NER, RE, QA, NLI, classification,
MIMIC-III sentence order prediction sentence similarity estimation
Clinical-Longformer L 149M Longformer MIMIC-III MLM NER, QA, NLI, classification
     Li et al. (2022a) 110M BigBird
BioBART Yuan et al. (2022a) L 140M, 406M BART PubMed sequence to sequence NER, entity linking,
summarization, dialogue
BioGPT Luo et al. (2022) L 355M, 1.5B GPT-2 PubMed next token prediction RE, QA, classification, generation
Med-PaLM Singhal et al. (2023a) L 8B, 62B, PaLM instructions instruction tuning QA
540B
GatorTron Yang et al. (2022b) L 345M, 3.9B, BERT Wikipedia, PubMed, MLM NER, RE, QA, NLI,
8.9B PMC, MIMIC-III, sentence similarity estimation
clinical narratives
ChatDoctor Li et al. (2023e) L 7B LLaMA HealthCareMagic instruction tuning dialogue
DoctorGLM Xiong et al. (2023) L 6B ChatGLM medical dialogues instruction tuning dialogue
BenTsao Wang et al. (2023d) L 7B LLaMA instructions instruction tuning QA, dialogue
MedAlpaca Han et al. (2023) L 7B, 13B LLaMA medical flash cards, instruction tuning QA
Stack Exchange,
WikiDoc
PMC-LLaMA Wu et al. (2024) L 7B, 13B LLaMA biomedical papers, next token prediction, QA
books, instructions instruction tuning
Med-PaLM 2 Singhal et al. (2023b) L 8B, 62B, PaLM 2 instructions instruction tuning QA
540B
HuatuoGPT Zhang et al. (2023b) L 7B, 13B BLOOM instructions instruction tuning QA, dialogue
MedCPT Jin et al. (2023c) L 110M BERT PubMed search logs retrieval classification, link prediction,
recommendation, retrieval,
sentence similarity estimation
Zhongjing Yang et al. (2024b) L 13B Ziya-LLaMA textbooks, QA pairs, next token prediction, QA
knowledge bases, EHRs, instruction tuning
EMRs, clinical reports,
instructions
DISC-MedLLM Bao et al. (2023) L 13B Baichuan instructions instruction tuning QA, dialogue
DRG-LLaMA Wang et al. (2024a) L 7B, 13B LLaMA MIMIC-IV classification diagnosis-related group prediction
Qilin-Med Ye et al. (2023b) L 7B Baichuan ChiMed-CPT, next token prediction, QA, dialogue
ChiMed-SFT, instruction tuning
ChiMed-DPO
AlpaCare Zhang et al. (2023d) L 7B, 13B LLaMA MedInstruct-52k instruction tuning QA, summarization
7B, 13B LLaMA-2
BianQue Chen et al. (2023d) L 6B ChatGLM BianQueCorpus instruction tuning dialogue
HuatuoGPT-II Chen et al. (2023a) L 7B, 13B, Baichuan 2 instructions instruction tuning QA, dialogue
34B
Taiyi Luo et al. (2024) L 7B Qwen instructions instruction tuning NER, RE, QA, classification
MEDITRON Chen et al. (2023e) L 7B, 70B LLaMA-2 GAP-Replay next token prediction, QA
instruction tuning
PLLaMa Yang et al. (2024c) L 7B, 13B LLaMA-2 plant science journals, next token prediction, QA
instructions instruction tuning
BioMistral Labrak et al. (2024) L 7B Mistral PMC next token prediction QA
Me LLaMA Xie et al. (2024) L 13B, 70B LLaMA-2 PubMed, PMC, next token prediction, NER, RE, QA, NLI, classification,
MIMIC-III, MIMIC-IV, instruction tuning summarization
MIMIC-CXR,
RedPajama, instructions
BiMediX Pieri et al. (2024) L 8$\times$7B Mixtral BiMed1.3M instruction tuning QA
MMedLM Qiu et al. (2024b) L 7B InternLM MMedC next token prediction QA
1.8B, 7B InternLM2
8B LLaMA-3
[…]
Hippocrates Acikgoz et al. (2024) L 7B LLaMA-2 PubMed, PMC, next token prediction, QA
7B Mistral medical guidelines, instruction tuning
instructions
BMRetriever Xu et al. (2024) L 410M, 1B Pythia biomedical papers, contrastive learning, QA, recommendation, retrieval,
2B Gemma textbooks, QA pairs, instruction tuning entity linking,
7B Mistral instructions sentence similarity estimation
Panacea Lin et al. (2024a) L 7B Mistral TrialAlign, next token prediction, summarization, query generation,
TrialInstruct instruction tuning query expansion, trial design,
patient-trial matching
G-BERT Shang et al. (2019) L+G BERT + MIMIC-III, ICD-9, MLM, diagnosis prediction, medication recommendation
GAT ATC medication prediction
CODER Yuan et al. (2022b) L+G 110M BERT UMLS link prediction entity linking, link prediction,
entity similarity estimation
MoP Meng et al. (2021b) L+G Adapters UMLS link prediction QA, NLI, classification
BioLinkBERT L+G 110M, 340M BERT PubMed MLM, NER, RE, QA, classification,
     Yasunaga et al. (2022b) link prediction sentence similarity estimation
[…]
(Biology and Medicine, Table Continued)
Model Modality Size Architecture Pre-training Data Pre-training Task(s) Evaluation Task(s)
ConVIRT Zhang et al. (2022) L+V BERT + MIMIC-CXR, text-image matching classification,
ResNet musculoskeletal text-to-image retrieval,
text-image pairs image-to-image retrieval
MMBERT Khare et al. (2021) L+V BERT + ROCO MLM VQA
ResNet
MedViLL Moon et al. (2022) L+V BERT + MIMIC-CXR MLM, VQA, classification,
ResNet text-image matching text-to-image retrieval,
image-to-text retrieval,
report generation
GLoRIA Huang et al. (2021) L+V BERT + CheXpert text-image matching classification, segmentation,
ResNet image-to-text retrieval
LoVT Müller et al. (2022) L+V BERT + MIMIC-CXR text-image matching segmentation, detection
ResNet
BioViL Boecking et al. (2022) L+V BERT + MIMIC-CXR MLM, NLI, classification, segmentation,
ResNet text-image matching phrase grounding
M3AE Chen et al. (2022c) L+V RoBERTa + ROCO, MedICaT MLM, VQA, classification,
ViT masked image modeling, text-to-image retrieval,
text-image matching image-to-text retrieval
ARL Chen et al. (2022d) L+V BERT + ROCO, MedICaT, MLM, VQA, classification,
ViT MIMIC-CXR masked image modeling, text-to-image retrieval,
text-image matching image-to-text retrieval
CheXzero Tiu et al. (2022) L+V Transformer + MIMIC-CXR text-image matching classification
ViT
MGCA Wang et al. (2022a) L+V BERT + MIMIC-CXR text-image matching classification, segmentation,
ResNet / ViT detection
MedCLIP Wang et al. (2022b) L+V BERT + MIMIC-CXR, text-image matching classification,
Swin CheXpert image-to-text retrieval
BioViL-T Bannur et al. (2023) L+V BERT + MIMIC-CXR MLM, classification, report generation,
ResNet text-image matching sentence similarity estimation
BiomedCLIP Zhang et al. (2023c) L+V BERT + PMC figure-caption text-image matching VQA, classification,
ViT pairs, fine-grained text-to-image retrieval,
text-image pairs image-to-text retrieval
PMC-CLIP Lin et al. (2023a) L+V BERT + PMC figure-caption MLM, VQA, classification,
ResNet pairs, subfigure- text-image matching text-to-image retrieval,
subcaption pairs image-to-text retrieval
Xplainer Pellegrini et al. (2023) L+V BERT + MIMIC-CXR text-image matching classification
ResNet
RGRG Tanida et al. (2023) L+V GPT-2 + MIMIC-CXR detection, classification, report generation
ResNet next token prediction
[…]
182M ResNet + PathVQA, Peir Gross, masked image modeling, summarization, image captioning,
[…]
OIA-DDR, CheXpert, VQA, treatment suggestion,
CytoImageNet, ISIC, image captioning mortality prediction
Retinal Fundus,
MIMIC-III, BioNLP,
PubMed
Med-UniC Wan et al. (2023) L+V BERT + MIMIC-CXR, text-image matching, classification, segmentation,
ResNet / ViT PadChest contrastive learning detection
LLaVA-Med Li et al. (2023a) L+V 7B LLaVA PMC figure-caption text-image matching, VQA
pairs, instructions instruction tuning
MI-Zero Lu et al. (2023a) L+V BERT + histopathology figure- text-image matching classification
CTransPath caption pairs
XrayGPT Thawkar et al. (2024) L+V LLaMA + MIMIC-CXR, text-image matching VQA
Swin Open-i
MONET Kim et al. (2024) L+V BERT + PMC and textbook text-image matching classification, data auditing,
ViT figure-caption pairs model auditing
QuiltNet Ikezogwo et al. (2023) L+V BERT + Quilt-1M text-image matching classification,
ViT text-to-image retrieval,
image-to-text retrieval
MUMC Li et al. (2023c) L+V BERT + ROCO, MedICaT, MLM, VQA
ViT ImageCLEFmedical text-image matching
Caption
M-FLAG Liu et al. (2023a) L+V BERT + MIMIC-CXR text-image matching classification, segmentation,
ResNet detection
PRIOR Cheng et al. (2023) L+V BERT + MIMIC-CXR text-image matching, classification, segmentation,
ResNet image reconstruction, detection,
sentence prototype generation image-to-text retrieval
Med-PaLM M Tu et al. (2024) L+V 12B, 84B, PaLM-E MultiMedBench instruction tuning QA, VQA, classification,
562B report generation,
report summarization
CITE Zhang et al. (2023i) L+V BERT + PatchGastric text-image matching, classification
ViT prompt tuning
Med-Flamingo Moor et al. (2023) L+V Flamingo PMC figure-caption next token prediction VQA
pairs, textbooks
RadFM Wu et al. (2023) L+V 14B LLaMA + MedMD, RadMD next token prediction, VQA, classification,
ViT instruction tuning report generation
PLIP Huang et al. (2023b) L+V GPT-2 + Twitter text-image pairs, text-image matching classification,
ViT PathLAION text-to-image retrieval,
image-to-image retrieval
MaCo Huang et al. (2023a) L+V BERT + MIMIC-CXR masked image modeling, classification, segmentation,
ViT text-image matching phrase grounding
CXR-CLIP You et al. (2023) L+V BERT + MIMIC-CXR, text-image matching, classification,
ResNet / Swin CheXpert, ChestX-ray14 contrastive learning image-to-text retrieval
Qilin-Med-VL Liu et al. (2023b) L+V LLaMA-2 + ChiMed-VL-Alignment, text-image matching, VQA
ViT ChiMed-VL-Instruction instruction tuning
BioCLIP Stevens et al. (2024) L+V GPT-2 + TreeOfLife-10M text-image matching classification
ViT
M3D Bai et al. (2024) L+V LLaMA-2 + M3D-Cap, M3D-VQA, text-image matching, VQA, segmentation,
ViT M3D-RefSeg, M3D-Seg instruction tuning text-to-image retrieval,
image-to-text retrieval,
report generation, 3D positioning
Med-Gemini Saab et al. (2024) L+V Gemini MedQA, LiveQA, instruction tuning QA, VQA, signal QA, video QA,
HealthSearchQA, classification,
MedicationQA, long-form text generation,
MIMIC-III, SLAKE, long EHR understanding
PathVQA, ROCO,
PAD-UFES-20,
MIMIC-CXR, ECG-QA
(Biology and Medicine, Table Continued)
Model Modality Size Architecture Pre-training Data Pre-training Task(s) Evaluation Task(s)
Med-Gemini-2D/3D/Polygenic L+V Gemini SLAKE, MIMIC-CXR, VQA, captioning, VQA, classification,
     Yang et al. (2024a) Digital Knee X-ray, instruction tuning report generation,
CXR-US2, NLST, disease risk prediction
CT-US1, PathVQA,
Histopathology,
PAD-UFES-20,
EyePACS, PMC-OA,
VQA-Med, UK Biobank
Mammo-CLIP Ghosh et al. (2024) L+V BERT + UPMC, VinDr-Mammo text-image matching classification, localization
EfficientNet
[…]
3B, 11B T5
[…]
contact prediction,
remote homology detection
[…]
contact prediction
[…]
     Chowdhury et al. (2022) chunk permutation prediction contact prediction
[…]
Gene Ontology remote homology detection,
fitness prediction
ProtGPT2 Ferruz et al. (2022) Protein 738M GPT-2 UniRef50 next token prediction secondary structure prediction,
disorder prediction,
protein sequence generation
ESM-IF1 Hsu et al. (2022) Protein 142M Transformer + UniRef50 next token prediction fixed backbone protein design,
GVP-GNN mutation effect prediction
ProGen Madani et al. (2023) Protein 1.6B CTRL UniParc, UniprotKB, next token prediction protein sequence generation
Pfam, NCBI Taxonomy
[…]
2.7B, 6.4B fitness prediction
[…]
150M, 650M, contact prediction,
3B, 15B 3D structure prediction
[…]
contact prediction,
embedding-based annotation
transfer,
remote homology detection,
fitness prediction,
localization prediction
[…]
text-protein matching localization prediction,
function annotation
[…]
ProtMPNN
[…]
text-guided protein editing,
secondary structure prediction,
contact prediction,
remote homology detection,
fitness prediction
[…]
398M, 898M R-GCN +
[…]
BioMedGPT Luo et al. (2023c) Protein 10B LLaMA-2 + S2ORC, PubChemQA, next token prediction, QA
GraphMVP + UniProtQA instruction tuning
ESM-2
[…]
fitness prediction,
localization prediction,
function annotation,
PPI prediction
BioT5 Pei et al. (2023) Protein 220M T5 C4, ZINC, UniRef50, sequence to sequence molecule property prediction,
PubMed, PubChem, protein property prediction,
Swiss-Prot drug-target interaction prediction,
PPI prediction,
molecule captioning,
text-based molecule design
ProLLaMA Lv et al. (2024) Protein 7B LLaMA-2 UniRef50, instructions next token prediction, protein sequence generation,
instruction tuning protein property prediction
DNABERT Ji et al. (2021) DNA 110M BERT GRCh38 MLM chromatin profile prediction,
promoter prediction,
splice site prediction,
functional genetic variant
identification
[…]
2.5B, 25B sequences prediction
[…]
     Dalla-Torre et al. (2023) 250M, 500M 1000 Genomes, enhancer prediction,
multispecies genomes promoter prediction,
epigenetic marks prediction,
splice site prediction
GENA-LM Fishman et al. (2023) DNA 110M, 340M BERT T2T-CHM13, MLM enhancer prediction,
110M BigBird 1000 Genomes, promoter prediction,
multispecies genomes epigenetic marks prediction,
splice site prediction,
species classification
DNABERT-2 Zhou et al. (2024a) DNA 110M BERT GRCh38, MLM chromatin profile prediction,
multispecies genomes promoter prediction,
epigenetic marks prediction,
splice site prediction,
species classification,
SARS-CoV-2 variant prediction,
enhancer-promoter interaction
(Biology and Medicine, Table Continued)
Model Modality Size Architecture Pre-training Data Pre-training Task(s) Evaluation Task(s)
HyenaDNA Nguyen et al. (2023a) DNA 0.4M, 3.3M, Hyena GRCh38 next token prediction chromatin profile prediction,
6.6M enhancer prediction,
promoter prediction,
epigenetic marks prediction,
splice site prediction,
species classification
[…]
6.6M sequence order prediction, chromatin profile prediction,
regression promoter prediction,
genomic signals and regions
recognition
[…]
     Akiyama and Sakakibara (2022) RNA clustering
[…]
3D structure prediction,
protein-RNA interaction,
mean ribosome load prediction
[…]
splice site prediction
[…]
solvent accessibility prediction
[…]
homologous sequences
prediction
[…]
classification, mRNA property prediction,
regression internal ribosome entry site
prediction
scBERT Yang et al. (2022a) Multi Performer PanglaoDB MLM cell type annotation,
novel cell type discovery
[…]
perturbation response prediction,
multi-batch integration,
multi-omic integration,
gene network inference
scFoundation Hao et al. (2024) Multi 100M Transformer + scRNA-seq data MLM cell clustering,
Performer drug response prediction,
perturbation response prediction,
cell type annotation,
gene network inference
[…]
chromatin dynamics prediction,
network dynamics prediction
CellLM Zhao et al. (2023b) Multi Performer PanglaoDB, MLM, classification, cell type annotation,
CancerSCEM contrastive learning drug sensitivity prediction
CellPLM Wen et al. (2024) Multi 82M Transformer scRNA-seq data, MLM cell clustering,
spatially-resolved scRNA-seq denoising,
transcriptomic data spatial transcriptomic imputation,
cell type annotation
Table A6: Summary of LLMs in geography, geology, and environmental science. “Climate”: Climate Time Series; “POI”: point of interest. Other notations have the same meaning as in previous tables.
Model Modality Size Architecture Pre-training Data Pre-training Task(s) Evaluation Task(s)
ClimateBERT L 82M DistilRoBERTa climate-related news, papers, MLM classification, fact-checking
       Webersinke et al. (2021) corporate climate reports
SpaBERT Li et al. (2022b) L 110M, 340M BERT OpenStreetMap MLM, entity typing, entity linking
masked entity prediction
[…]
masked geographic modeling,
contrastive learning
K2 Deng et al. (2024) L 7B LLaMA geoscience papers, next token prediction, QA
Wikipedia, instructions instruction tuning
OceanGPT Bi et al. (2023b) L 7B LLaMA-2 ocean science papers, next token prediction, QA, classification, extraction,
instructions instruction tuning knowledge probing,
commonsense reasoning,
summarization, generation
ClimateBERT-NetZero L 82M DistilRoBERTa Net Zero Tracker classification classification
       Schimanski et al. (2023)
GeoLM Li et al. (2023f) L 110M, 340M BERT OpenStreetMap, MLM, NER, RE, entity typing,
Wikipedia contrastive learning entity linking
GeoGalactica Lin et al. (2024c) L 30B Galactica geoscience papers, code, next token prediction, QA, knowledge probing,
Wikipedia, instructions instruction tuning quantitative reasoning,
summarization, generation
ERNIE-GeoL Huang et al. (2022) L+G Transformer + Baidu Maps MLM, classification,
graph aggregation (POI database, geocoding query-POI matching,
search logs) address parsing, geocoding,
next POI recommendation
[…]
Knowledge Graph bag-of-words prediction,
classification
UrbanCLIP Yan et al. (2024) L+V Transformer + satellite images, next token prediction, urban indicator prediction
ViT location descriptions, text-image matching
[…]
climate projection,
climate model downscaling
FengWu Chen et al. (2023b) Climate Transformer ERA5 regression weather forecasting
W-MAE Man et al. (2023) Climate ViT ERA5 masked image modeling weather forecasting
[…]